Which of the following trinomials is prime over the set of rational numbers?
A x2 – 9x + 12
B x2 – x – 12
C x2 + 4x – 12
D x2 + 7x + 12

Answer by ewatrrr(24785)


Hi
Note: trinomials that are prime over the set of rational numbers will NOT factor.
A Tinomial ax^2 + bx + c 'will factor' if (b^2-4ac)is a perfect square or zero.
A. x2 – 9x + 12 81-48 = 33 Prime ( 33 is NOT a perfect square)
B. x2 – x – 12 1+48 = 49 (49 is a perfect square) factoring results: (x-4)(x+3)
C. x2 + 4x – 12 16+48 = 64 (64 is a perfect square) factoring results: (x+6)(x-2)
D. x2 + 7x + 12 49-48 = 1 (1 is a perfect square) factoring results: (x+3)(x+4)